OpenStudy (bohotness):

help please will medal and fan 15 question

OpenStudy (shaadykhamlichi):

?????/

OpenStudy (anonymous):

name one just one at a time though, otherwise gets too complicated

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (bohotness):

okay

OpenStudy (anonymous):

@satellite73 how did you get to be smart score 100

OpenStudy (bohotness):

What is the volume of a cube, if the length of one side is 9 inches? A. 486 cubic inches B. 27 cubic inches C. 81 cubic inches D. 729 cubic inches ci think ?

OpenStudy (shaadykhamlichi):

i think d

OpenStudy (anonymous):

volume of a cube is the side cubed (hence the name) compute \(9^3\)

OpenStudy (shaadykhamlichi):

because volume of a cube = a^3=a*a*a

OpenStudy (anonymous):

it is \(9^3=729\) not \(9^2=81\) since it is a CUBE and not a SQUARE

OpenStudy (shaadykhamlichi):

yup

OpenStudy (shaadykhamlichi):

that's true 100%